Sandhill Chemicals Company acquires a delivery truck at a cost of $40,000 on January 1, 2022. The truck is expected to have a salvage value of $3,400 at the end of its 5-year useful life. Assuming the declining-balance depreciation rate is double the straight-line rate, compute annual depreciation for the first and second years under the declining-balance method.

First Year

Second Year

Annual depreciation under declining-balance method

$Enter a dollar amount

$Enter a dollar amount

eTextbook and Media
Answer #1

Double declining rate = 200/Useful life = 200/5 = 40%

First year

= 40,000*40%

= 16,000

Second Year

= (40,000-16,000)*40%

= 9,600